OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 4 − (x + 2) = 6 − 5(x + 4) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (sh3lsh):

In your words, not ours. Be able to solve and you'll know the answer quickly!

OpenStudy (anonymous):

Okay well distribute the negative in front of the (x+2) and the -5 in front of the (x+4) so it looks like this: 4-x-2= 6-5x+20 now you can simplify that into: 2-x=26-5x now add 5x to both sides, and subtract 2 from both sides. You'd get 4x=24 now divide both sides by 4... x=6 :)

OpenStudy (anonymous):

@shelah yhu gptt 2.0 secconds to get off mhyy stuff i ant fa no disagreement